1. In an experiment to study the relationship of hypertension and smoking habits, the following data below are collected for 180 individuals, where H and NH in the table stands for Hypertension and Non-hypertension, respectively. If one of these individuals is selected at random, find the probability that the person is:

(a) not experiencing hypertension, given that the person is a heavy smoker
(b) a nonsmoker, given that the person is not experiencing hypertension
(c) either a moderate or heavy smoker, given that the person is not experiencing hypertension
1
Expert's answer
2020-11-03T11:54:19-0500

We have the following dara


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c }
 & Nonsmokers & Moderate\ & Heavy & Total \\ 
 & & smokers & smokers & \\ \hline
 H & 21 & 36 & 30 & 87 \\ 
 NH & 48 & 26 & 19 & 93 \\ 
 \hdashline
 Total & 69 & 62 &49 & 180
\end{array}"

where H and NH stands for Hypertension and NonHypertension, respectively.

Let "N_S" denotes the nonsmoker, "M_S" denotes the moderate smoker, and "H_S" denotes the heavy smoker.

One of these inividuals is selected at random.

a)

"P(NH|H_S)=\dfrac{19}{49}"

b)

"P(N_S|NH)=\dfrac{48}{93}=\dfrac{16}{31}"

c)


"P((M_S\cup H_S)|NH)=\dfrac{26+19}{93}=\dfrac{15}{31}"